### v3

# Foundations of Reasoning in Language Models (FoRLM)

Abstract
Test-time reasoning architectures such as those following the Generate-Verify paradigm, where a model iteratively refines or verifies its own generated outputs, prioritise generation and verification but exclude the monitoring processes that determine when and how reasoning should begin. This omission may contribute to the prefix dominance trap, in which models commit early to suboptimal reasoning paths and seldom recover, yielding roughly 20% accuracy loss. We address this architectural gap by proposing the Monitor-Generate-Verify (MGV) framework, a computational translation of Flavell’s and Nelson and Narens’ metacognitive theories that preserves their psychological detail. MGV extends the Generate-Verify paradigm by adding explicit monitoring that captures metacognitive experiences (from difficulty assessments to confidence judgements) before generation begins and refines future monitoring through verification feedback. Though we present no empirical validation, MGV provides a vocabulary for diagnosing component-level failures in reasoning systems, suggests specific architectural interventions for future designs, and identifies connections to resource-rational analysis that may ground its mechanisms in normative principles.
1 Introduction
Once language models commit to an initial reasoning strategy, subsequent verification rarely helps; this prefix dominance trap causes nearly 20% performance degradation when models choose suboptimal approaches, with limited recovery possible through refinement (Luo et al., 2025). Today’s dominant Generate-Verify test-time reasoning architectures (Weng et al., 2023; Madaan et al., 2023; Lee et al., 2025; Zhang et al., 2024) exemplify this limitation through their very design. They operate through immediate generation followed by iterative refinement, without assessing task characteristics or selecting appropriate strategies before generating solutions. What architectural vocabulary might help diagnose what these systems lack?
Cognitive psychology offers a candidate answer. Before attempting complex tasks, humans engage in metacognitive monitoring, assessing difficulty, retrieving relevant strategies, and establishing confidence criteria, often without conscious deliberation (Flavell, 1979; Nelson and Narens, 1990). This monitoring capacity transforms uncertainty into tractable signals that guide subsequent action. Flavell’s model specifies how metacognitive knowledge and experience interact to enable strategy selection, while Nelson and Narens’ metamemory framework details how confidence thresholds and adaptive search mechanisms emerge from hierarchical monitoring and control. These theories suggest that the Generate-Verify paradigm may be missing an entire phase of processing that precedes generation.
However, translating these psychological insights into architectural recommendations requires computational specification. Resource-rational analysis has made significant progress here, with Callaway et al. (2024) formalising Nelson and Narens’ metamemory theory as a meta-level MDP that derives optimal stopping policies from cost-benefit principles. While their approach offers normative grounding and empirical validation, the present work pursues a complementary goal: instead of deriving optimal behaviour from computational constraints, we translate the detailed psychological vocabulary of Flavell (1979) and Nelson and Narens (1990) into algorithmic form, preserving structure that resource-rational approaches may abstract away.
The resulting Monitor-Generate-Verify (MGV) framework extends the Generate-Verify paradigm by adding explicit monitoring that captures metacognitive experiences before generation begins and refines future monitoring through verification feedback. Our formalisation preserves constructs including the tripartite organisation of metacognitive knowledge, the distinction between difficulty and evaluative experiences, dual-counter evidence accumulation for feeling-of-knowing, and satisficing threshold dynamics. We offer no empirical validation and no normative justification for these architectural choices. The contribution is a detailed computational translation that provides diagnostic vocabulary for understanding what reasoning systems might lack, suggesting where architectural interventions might prove productive.
The remainder of this paper is organised as follows. Section 2 situates our contribution within resource-rational approaches to metacognition and LLMs. Section 3 presents algorithmic formalisations of Flavell’s and Nelson and Narens’ theories, translating psychological concepts into computational structures. Section 4 acknowledges limitations and outlines future directions.
2 Related Work
This section situates the present work within two related research programmes. The first is resource-rational analysis, a cognitive modelling paradigm that formalises metacognition as cost-benefit optimisation over computational resources. The second is recent work applying these principles to language models. We review each in turn before clarifying how the present work differs in its foundational commitments.
2.1 Resource-Rational Analysis of Metacognition
Resource-rational analysis is a cognitive modelling paradigm that reconceives bounded rationality not as a departure from optimality but as optimality under constraints (Griffiths et al., 2015, 2019; Lieder and Griffiths, 2020). Making good decisions requires computation, yet computation itself is costly. This tension between the value of deliberation and its cost gives rise to a class of meta-problems: problems not about what to decide, but about how to decide. How much should an agent think before acting? Which cognitive strategy should be deployed for a given problem? Which aspects of a decision path merit exploration, and which can be safely ignored? These questions concern the allocation of computational resources, and answering them requires reasoning about reasoning itself.
Resource-rational analysis provides a framework for addressing such meta-problems by treating cognitive constraints as part of the optimisation problem rather than as departures from rationality. Under metalevel rationality, an agent is assessed not by the expected utility of actions taken, but by how well its cognitive algorithm trades off decision quality against deliberation costs. From this perspective, rationality concerns not only making good decisions but also employing efficient cognitive strategies. The framework reduces meta-reasoning [^1] Meta-reasoning should be distinguished from meta-learning. Meta-reasoning concerns how to allocate limited computational resources during cognition, deciding which computations to perform and when to stop deliberating. Meta-learning concerns how to structure the learning process itself, exploiting knowledge about the statistical structure of learning environments to make better use of limited data. Both address resource limitations, but at different levels: meta-reasoning optimises the use of time and cognitive effort within a task, whereas meta-learning optimises the acquisition of knowledge across tasks. For a treatment that situates both perspectives within the resource-rational framework, see Griffiths et al. (2019). to a decision-theoretic problem: for any computation the agent could perform, there exists a value of computation (VOC) that quantifies the expected improvement in decision quality minus the cost of deliberation.
This cognitive modelling paradigm provides a common structure that different research programmes have instantiated in distinct ways, each addressing a different facet of the meta-reasoning problem. Lieder and Griffiths (2017) modelled strategy selection: rather than asking whether to think more, this formulation asks how to think, treating the choice among cognitive strategies as a contextual bandit in which the agent learns to predict which mode of reasoning best suits each type of problem (Appendix 7.2). Lieder et al. (2018) modelled cognitive control: by casting control specification as a belief-MDP and introducing a learned approximation, this formulation explains how control exhibits plasticity, adapting through experience to select signals that optimally trade off effort against goal achievement (Appendix 7.4). Callaway et al. (2022) modelled planning: by embedding deliberation within a meta-level MDP, this formulation captures the sequential structure of mental simulation, recognising that the value of a computation depends not only on its immediate effect but also on the future computations it enables (Appendix 7.3). Callaway et al. (2024) modelled memory recall: by treating retrieval as evidence accumulation supervised by an optimal stopping process, this formulation reveals the adaptive function of feeling-of-knowing, providing the monitoring signal that enables efficient termination of searches unlikely to succeed while sustaining effort when retrieval remains probable (Appendix 7.5).
These four instantiations span a range of formal structures, from contextual bandits to belief-MDPs, and a range of cognitive phenomena, from high-level strategy choice to low-level memory retrieval. Yet they share a common theoretical commitment: the mind’s computational processes are themselves subject to rational optimisation. Appendix 7 provides a detailed technical treatment of these frameworks, comparing their formal structures and identifying shared algorithmic motifs. Here, we turn to recent work that has begun extending these principles beyond cognitive modelling to the training of language models.
2.2 Rational Metareasoning for Language Models
While the research reviewed above models human cognition, recent work has begun applying the VOC framework directly to language model training. De Sabbata et al. (2024) introduced RaM (Rational Metareasoning), which trains LLMs to use chain-of-thought reasoning selectively based on task difficulty. Their approach defines the reward of a reasoning chain $z$ as the difference between its utility and cost:
$R_{\pi}(x,y,z)=U_{\pi}(z|x,y)-C(z),$
where utility $U_{\pi}(z|x,y)=\log\pi_{\theta}(y|z,x)-\log\pi_{\theta}(y|x)$ quantifies the increase in likelihood of generating the correct answer $y$ when the chain of thought $z$ is included, and cost $C(z)=\gamma\cdot l(z)$ is proportional to the number of tokens. This formulation mirrors the equation, with reasoning tokens corresponding to intermediate computations and model outputs corresponding to actions. Models trained with this objective learn to generate shorter reasoning chains for easier problems and longer chains for harder problems, demonstrating that VOC-based training can induce adaptive computation in neural systems. This work provides empirical evidence that resource-rational principles can be operationalised in modern architectures, though it focuses on optimising existing reasoning capabilities rather than introducing new metacognitive components.
2.3 Relationship to the Present Work
Formalising metacognitive theory requires choosing what to treat as primitive and what to derive. Resource-rational analysis treats computational operations as primitive and derives metacognitive behaviour from cost-benefit optimisation. The present work makes the opposite choice, treating psychological constructs as primitive and translating them directly into algorithmic form. Put simply, instead of deriving that people think the way they do because it is efficient given limited cognitive resources, this work treats psychological ideas as the foundation and directly formalises them in computational terms.
This difference in approach leads to different contributions. Resource-rational analysis provides normative grounding, explaining why certain metacognitive policies are optimal given computational constraints. The present work provides psychological granularity, preserving detailed structure from the source theories that resource-rational models abstract away. Section 3 makes this concrete by formalising constructs including tripartite metacognitive knowledge, dual-counter evidence accumulation, and satisficing threshold dynamics. Section 4 returns to the relationship between these approaches and examines how they might complement one another.
3 Monitor-Generate-Verify (MGV)
Flavell (1979) and Nelson and Narens (1990) developed seminal theories of how metacognition coordinates cognitive processes through monitoring and control loops. These frameworks, though developed for human cognition, offer potential blueprints for computational systems. Flavell’s model provides a dynamic architecture where metacognitive knowledge and experience guide strategy selection and verification, while Nelson and Narens’ metamemory framework specifies how confidence thresholds and adaptive search mechanisms emerge from hierarchical monitoring and control. By computationally formalising these psychological theories, we establish Monitor-Generate-Verify (MGV) as a theoretical framework for understanding how explicit metacognitive mechanisms could address the architectural limitations of current reasoning systems. The following subsections present detailed formalisations that translate these cognitive psychology insights into algorithmic structures, revealing both what current architectures lack and how metacognitive principles might be operationalised computationally.
3.1 Flavell’s Model of Metacognition
Flavell (1979) conceptualises metacognition as a dynamic control architecture comprising four interacting components: metacognitive knowledge, metacognitive experience, goals (or tasks), and actions (or strategies). Rather than operating as independent modules, these components form an integrated system characterised by continuous bidirectional influences, positioning metacognition as a self-regulating system capable of adaptive control over cognitive processes. We present the core computational structure below, with a complete mathematical formalisation provided in Appendix 5.
3.1.1 Cognitive Monitoring
The regulation process begins with initialisation, where task $\mathcal{T}$ and goal $\mathcal{G}$ establish the initial state $\mathcal{S}_{0}=f(\mathcal{T},\mathcal{G})$. While Flavell (1979) treats goals and tasks as equivalent, we maintain a computational distinction. $\mathcal{T}$ represents the cognitive enterprise while $\mathcal{G}$ specifies success criteria, enabling clearer analysis of metacognitive processes.
Algorithm 1 Flavell’s Model of Cognitive Control
1:Initialise: $\mathcal{S}_{0}\leftarrow f(\mathcal{T},\mathcal{G})$; $\tau\leftarrow 0$
2:while $\mathcal{S}_{\tau}=\text{ACTIVE}$ do
3:  // MONITOR: Retrieve knowledge & assess experience
4:  $\mathcal{MK}_{\tau}\leftarrow$ if $\tau=0$ then retrieve($\mathcal{MK},\mathcal{T},\mathcal{G}$)
5:      else $\mathcal{MK}_{\tau-1}\cup$ retrieve($\mathcal{MK},\mathcal{ME}_{\tau-1}$)
6:  $\mathcal{ME}_{\text{difficulty}}^{\tau}\leftarrow$ feel($\mathcal{T}$, Outcomesτ-1) $\oplus$ assess($\mathcal{T}$, $\mathcal{MK}_{\tau}$)
7:  // GENERATE: Select & execute cognitive strategy
8:  $\mathcal{CS}_{\tau}\leftarrow$ select($s\in\mathcal{MK}_{\text{Strategy}}\mid\mathcal{ME}_{\text{difficulty}}^{\tau},\mathcal{MK}_{\tau},\mathcal{T},\mathcal{G}$)
9:  $\mathcal{CO}_{\tau}\leftarrow$ execute($\mathcal{CS}_{\tau},\mathcal{T},\mathcal{G}$)
10:  // VERIFY: Evaluate progress & update knowledge
11:  $\mathcal{ME}_{\text{evaluative}}^{\tau}\leftarrow$ assess($\mathcal{CO}_{\tau},\mathcal{MK}_{\tau}$)
12:  $\mathcal{MS}_{\tau}\leftarrow$ select($s\in\mathcal{MK}_{\text{Strategy}}^{\text{meta}}\mid\mathcal{ME}_{\text{evaluative}}^{\tau}$)
13:  $\mathcal{MO}_{\tau}\leftarrow$ execute($\mathcal{MS}_{\tau},\mathcal{CO}_{\tau},\mathcal{MK}_{\tau},\mathcal{G}$)
14:  $\mathcal{MK}\leftarrow$ update($\mathcal{MK}$, $\Phi_{\tau}$) where $\Phi_{\tau}=(\mathcal{ME}_{\tau},\text{Strategy}_{\tau},\text{Outcome}_{\tau})$
15:  $\mathcal{S}_{\tau+1}\leftarrow$ if goal_achieved($\mathcal{CO}_{\tau},\mathcal{G}$) then TERMINATE else ACTIVE
16:  $\tau\leftarrow\tau+1$
17:end while
The monitoring phase activates metacognitive knowledge differently across cycles. Initial cycles rely solely on task-goal combinations, while subsequent cycles incorporate emerging metacognitive experiences from $\tau-1$ that trigger additional relevant knowledge. According to Flavell (1979), this knowledge comprises three categories: agent variables ($\mathcal{MK}_{\text{Agent}}$) representing learned self-models of performance patterns and processing preferences; task variables ($\mathcal{MK}_{\text{Task}}$) capturing knowledge about cognitive situation assessment including information characteristics and task demands; and strategy variables ($\mathcal{MK}_{\text{Strategy}}$) encompassing knowledge about the effectiveness of both cognitive strategies (problem-solving procedures) and metacognitive strategies (monitoring and regulation processes). These categories function as an integrated system where task variables diagnose cognitive demands, strategy variables prescribe responses, and agent variables contextualise both within the agent’s capabilities.
The monitoring phase also generates metacognitive experiences of difficulty ($\mathcal{ME}_{\text{difficulty}}^{\tau}$), which Flavell (1979, p. 909) describes as subjective feeling-of-complexity, comprehension challenges, or sensing that material exceeds current capabilities. These experiences evolve through iterative assessments, progressing from initial coarse feelings to increasingly nuanced evaluations of specific challenge sources.
During the generation phase, metacognitive experiences function as computational signals that require interpretation through metacognitive knowledge to guide strategy selection. The process follows a two-phase pattern. First, $\mathcal{MK}_{\text{Strategy}}$ transforms general difficulty signals into precise diagnostic patterns (e.g., “content uncertainty with unknown terms” or “procedural confusion from missing steps”). Second, these refined patterns activate corresponding cognitive strategies. The selected strategy $\mathcal{CS}_{\tau}$ is then executed to produce cognitive outcomes $\mathcal{CO}_{\tau}$, generating feedback that provides both task progress information and context for subsequent monitoring.
The verification phase evaluates these outcomes, triggering what Flavell (1979, p. 909) describes as additional metacognitive experiences about performance rather than difficulty. These evaluative experiences ($\mathcal{ME}_{\text{evaluative}}^{\tau}$) activate metacognitive strategies that assess whether outcomes form a coherent whole, appear plausible and consistent with prior knowledge, and provide an avenue to the goal. The specific metacognitive strategy $\mathcal{MS}_{\tau}$ selected depends on the nature of the evaluative signal: uncertainty about validity triggers plausibility checking, sensing incompleteness activates coherence assessment, and so forth. Notably, these experiences can add to, delete from, or revise the metacognitive knowledge base through Piagetian mechanisms (Flavell, 1963), with the complete experience tuple $\Phi_{\tau}$ updating $\mathcal{MK}$ for future cycles.
3.1.2 Memory and Learning Gaps
A significant limitation in Flavell’s model is the absence of explicit working memory mechanisms for storing information across monitoring cycles. The model does not specify where $\mathcal{ME}_{\text{difficulty}}^{\tau}$ resides during strategy execution, how $\mathcal{CO}_{\tau}$ is maintained during evaluation, or how experience patterns across cycles are retained for subsequent processing. This absence precludes sophisticated termination criteria that would require access to historical monitoring data across the complete sequence $\Phi=(\Phi_{0},\Phi_{1},\ldots,\Phi_{T})$.
With an explicit memory component, the model could implement comprehensive abandonment criteria that evaluate: (1) repeated strategy failures indicated by consistently negative $\mathcal{MO}_{\tau}$ across multiple cycles, suggesting task intractability; (2) resource constraints where cumulative effort across $\Phi_{0}$ to $\Phi_{\tau}$ exceeds acceptable limits relative to $\mathcal{MK}_{\text{Agent}}$; (3) goal displacement where evolving $\mathcal{ME}_{\text{evaluative}}^{\tau}$ signals that alternative objectives have become more salient than the original $\mathcal{G}$; and (4) insurmountable goal-state discrepancy where the pattern of $\mathcal{CO}_{\tau}$ outcomes reveals fundamental incompatibility with $\mathcal{G}$ achievement.
A related temporal limitation concerns metacognitive knowledge acquisition and refinement. While Flavell acknowledges that experiences can ‘add to’, ‘delete from’, or ‘revise’ the knowledge base, the model assumes pre-existing $\mathcal{MK}$ without specifying learning mechanisms – how unsuccessful strategies refine strategy knowledge, or how repeated encounters improve task assessments.
Such memory-dependent termination decisions and learning-dependent knowledge refinement would better reflect real-world metacognitive monitoring, where individuals track cumulative progress patterns and recognise when persistence becomes counterproductive, while simultaneously refining their metacognitive knowledge through experience. These limitations point towards the necessity for more sophisticated architectural frameworks that explicitly model the temporal dynamics of metacognitive information storage and retrieval as well as the acquisition and refinement of metacognitive knowledge – considerations that become central to Nelson and Narens’ metamemory architecture.
3.2 Nelson and Narens’ Model of Metamemory
Nelson and Narens (1990) establish metacognition as fundamentally hierarchical, distinguishing between object-level and meta-level processes. The object-level comprises the cognitive operations themselves, such as searching memory for a target item. The meta-level supervises these operations, by monitoring progress toward the goal (e.g., how close is recall?) and controlling how long the process is allowed to continue (e.g., should the search persist or terminate?). This supervision relies on two distinct information flows: monitoring conveys information upward from object-level to meta-level, while control conveys decisions downward from meta-level to object-level. These relationships are logically independent and asymmetric: the meta-level maintains a model of the object-level, while the object-level operates without corresponding meta-level representation. We present the core computational structure below, with complete mathematical formalisation provided in Appendix 6.
3.2.1 Acquisition Process
The acquisition process begins with establishing the norm of study $\mathcal{N}_{s}=\rho^{*}\times(1+\delta_{\text{retention}})$, where $\rho^{*}$ represents target performance and $\delta_{\text{retention}}$ captures beliefs about memory decay over interval $\tau_{\text{delay}}$. This operationalises abstract goals into quantified mastery criteria that anticipate forgetting. Following Ericsson and Simon (1984), monitoring occurs within working memory (STM), with information from long-term memory (LTM) accessed probabilistically via $\text{retrieve}_{\theta}(\cdot)$ where $\theta$ represents access probability (Atkinson and Shiffrin, 1968).
Algorithm 2 Nelson and Narens’ Model of Acquisition
[…]
4:while $\mathcal{J}_{\tau}\neq\emptyset$ do
5:  // MONITOR: Assess mastery via EOL/FOK
[…]
7:  for each $j\in\mathcal{J}_{\tau}$ do
8:   $\mathcal{ME}_{\tau,j}[1]\leftarrow$ if $\tau=1$ then EOL($i_{j}$) else FOK($i_{j},\mathcal{CO}_{\tau-1,j}$)
9:  end for
10:  // GENERATE: Allocate resources & select strategies
11:  for each $j\in\mathcal{J}_{\tau}$ do
12:   $r_{\tau,j}\leftarrow R_{\text{total}}\times(\mathcal{ME}_{\tau,j}[1])^{-1}/\sum_{k}(\mathcal{ME}_{\tau,k}[1])^{-1}$
[…]
15:  end for
16:  // VERIFY: Judge learning & update items
17:  for each $j\in\mathcal{J}_{\tau}$ do
[…]
19:   $\mathcal{ME}_{\tau,j}[2]\leftarrow$ JOLτ,j
20:   $\Phi_{\tau}^{\text{STM}}\leftarrow\Phi_{\tau}^{\text{STM}}\cup\{(\mathcal{ME}_{\tau,j},i_{j},r_{\tau,j},\sigma_{\tau,j},\mathcal{CO}_{\tau,j})\}$
21:  end for
22:  $\mathcal{J}_{\tau+1}\leftarrow\{j\in\mathcal{J}_{\tau}:\mathcal{N}_{s}-\text{JOL}_{\tau,j}>0\}$; $\tau\leftarrow\tau+1$
23:end while
[…]
The monitoring phase generates metacognitive experiences as multidimensional vectors. Ease-of-learning (EOL) provides initial difficulty assessment, while feeling-of-knowing (FOK) incorporates prior outcomes to refine mastery judgements. These phenomenological experiences serve as primary input for control decisions (Nelson and Narens, 1990, p. 160). During the generation phase, resource allocation operates inversely to EOL/FOK values – items with lower metacognitive confidence receive proportionally more resources $r_{\tau,j}=R_{\text{total}}\times w_{j}/\sum_{k}w_{k}$ where $w_{j}=(\mathcal{ME}_{\tau,j}[1])^{-1}$. Strategy selection integrates these metacognitive inputs to map appropriate learning methods to individual items.
The verification phase employs judgement-of-learning (JOL) to evaluate mastery following cognitive outcomes. Items achieving the norm of study ($\text{JOL}_{\tau,j}\geq\mathcal{N}_{s}$) are removed from further consideration, while those below threshold remain in $\mathcal{J}_{\tau+1}$ for continued learning. The complete experience tuple accumulates in working memory as $\Phi_{\tau}^{\text{STM}}$, subsequently undergoing consolidation to LTM at encoding rate $\psi$.
3.2.2 Retrieval Process
The retrieval process implements Nelson and Narens’ dual-counter FOK hypothesis, where $\text{FOK}^{+}$ accumulates evidence for information presence whilst $\text{FOK}^{-}$ accumulates evidence for absence, consistent with ‘knowing not’ (Kolers and Palef, 1976). Initial thresholds are personalised through metacognitive calibration history: $\lambda_{\text{FOK}}^{(0)}=\text{median}(\{||\text{FOK}||:\text{successful retrievals in }\mathcal{MK}_{0}^{\text{STM}}\})$ and $\lambda_{\text{confidence}}^{(0)}=\text{median}(\{\text{confidence}:\text{correct outputs in }\mathcal{MK}_{0}^{\text{STM}}\})$, embodying the No-Magic Hypothesis by utilising recallable metacognitive knowledge.
Algorithm 3 Nelson and Narens Model of Retrieval
[…]
3:while search active do
4:  // MONITOR: Assess dual-counter FOK
[…]
7:  // Determine search intensity based on FOK evidence
8:  if $\text{FOK}_{\tau}||<\lambda_{\text{FOK}}^{(\tau)}$ then
9:   $\mathcal{S}_{\tau}\leftarrow\text{ACTIVE}_{\text{intensive}}$ $\triangleright$ Insufficient evidence
10:  else if $\text{FOK}_{\tau}^{+}>\text{FOK}_{\tau}^{-}$ then
11:   $\mathcal{S}_{\tau}\leftarrow\text{ACTIVE}_{\text{standard}}$ $\triangleright$ Positive dominance
12:  else
13:   break $\triangleright$ Negative dominance: terminate
14:  end if
15:  // GENERATE: Attend to cues & automatic search
[…]
18:  // VERIFY: Evaluate answer & adjust thresholds
[…]
21:   output $\mathcal{A}_{\tau}$; break
22:  else if $\mathcal{A}_{\tau}=$ null $\wedge\text{FOK}_{\tau}^{-}>\text{FOK}_{\tau}^{+}$ then
23:   output null; break $\triangleright$ Omission
24:  end if
25:  $\Omega_{\tau}^{\text{STM}}\leftarrow\Omega_{\tau}^{\text{STM}}\cup\{(\text{FOK}_{\tau},\text{cue}_{\tau},\mathcal{A}_{\tau},\text{confidence}_{\tau})\}$
26:  $\beta_{\tau}\leftarrow\exp(-\alpha\cdot(\tau+|\{$failed attempts in $\Omega_{\tau}^{\text{STM}}\}|))$
27:  $\lambda_{\text{confidence}}^{(\tau+1)},\lambda_{\text{FOK}}^{(\tau+1)}\leftarrow\lambda^{(0)}\cdot\beta_{\tau}$ $\triangleright$ Satisficing
28:  $\tau\leftarrow\tau+1$
29:end while
[…]
The monitoring phase employs rapid FOK assessment that operates faster than actual recall (Reder, 1987), enabling efficient search control. When FOK magnitude falls below threshold ($\text{FOK}_{\tau}||<\lambda_{\text{FOK}}^{(\tau)}$), insufficient evidence triggers intensive cue attention to gather additional metacognitive information. With sufficient evidence, positive dominance ($\text{FOK}_{\tau}^{+}>\text{FOK}_{\tau}^{-}$) warrants continued search, while negative dominance justifies termination.
The generation phase reflects Nelson and Narens’ insight that search execution is automatic once initiated – conscious control operates through cue attention intensity rather than strategy selection. The automatic search process $\text{search}_{\text{auto}}(\text{cue}_{\tau})$ operates through pattern recognition, potentially yielding identical results across consecutive cycles due to its deterministic nature.
Verification distinguishes two error pathways: commission errors (outputting incorrect answers with high confidence) and omission errors (terminating without answers following prolonged search). Following satisficing principles (Simon, 1979), both confidence and FOK thresholds undergo dynamic adjustment: $\lambda^{(\tau+1)}=\lambda^{(0)}\cdot\beta_{\tau}$ where $\beta_{\tau}=\exp(-\alpha\cdot\text{burden})$ captures accumulating search costs. This ensures previously inadequate answers may become acceptable as search burden increases, preventing exhaustive search behaviour.
3.2.3 Memory Consolidation and Knowledge Evolution
A distinctive strength of Nelson and Narens’ framework lies in its explicit treatment of long-term memory (LTM) as both a repository and an evolving knowledge base. During acquisition and retrieval, the experience tuples accumulated in working memory ($\Omega_{T}^{\text{STM}}$) undergoes consolidation into LTM at encoding rate $\psi$:
While Nelson and Narens do not explicitly specify the timing of this consolidation process, it likely occurs during the verification stage at rate $\psi$, potentially operating below conscious awareness. This consolidation mechanism enables the global metacognitive knowledge base to evolve through accumulated experience, distinguishing Nelson and Narens’ approach from more static metacognitive frameworks. The probabilistic retrieval function $\text{retrieve}_{\theta}(\mathcal{MK},\cdot)$ subsequently accesses this enriched knowledge base, creating a dynamic feedback loop where metacognitive experiences inform future metacognitive assessments.
4 Limitations and Future Work
4.1 From Specification to Implementation
The MGV framework specifies what should be computed but not how to compute it in neural architectures. Operationalising constructs like $\mathcal{ME}_{\text{difficulty}}$ or dual-counter FOK requires identifying appropriate neural correlates or designing explicit computational mechanisms. For language models, this might involve obtaining logit-based confidence ratings (Kumaran et al., 2025), using entropy-based proxies for metacognitive experiences, or implementing explicit evidence accumulators. Recent work suggests that LLMs access metacognitive signals they cannot adequately articulate, with implicit confidence measures outperforming explicit verbalisation (Wang et al., 2025; Tian et al., 2023; Xiong et al., 2023; Cash et al., 2024; Griot et al., 2025; Lindsey et al., 2025). Neural analysis reveals a lower-dimensional “metacognitive space” where monitoring signals correspond to linearly separable directions (Ji-An et al., 2025; Zou et al., 2023; Liu et al., 2023), consistent with cognitive findings that metacognition operates on abstracted representations (Reder, 1987).
4.2 Scope of Formalisation
The present work formalises two foundational theories, but metacognition research extends well beyond Flavell and Nelson and Narens. A more complete computational account of metacognition should integrate subsequent theoretical developments. Additionally, our formalisations make interpretive choices where the source theories are ambiguous or silent. For instance, the exponential decay of satisficing thresholds reflects our reading of the theories rather than explicit specifications in the original texts. Alternative formalisations are possible and might yield different architectural implications.
4.3 Normative Grounding
The framework lacks principled justification for its architectural choices. Resource-rational analysis explains why certain metacognitive policies are optimal given computational constraints. MGV, by contrast, specifies mechanisms without explaining why those mechanisms are appropriate. As Callaway et al. (2024, pp. 17–18) noted regarding related work in metamemory, this paper shares the limits of Metcalfe (1993), Koriat and Goldsmith (1996), and Bennett et al. (2017) in focusing on functional aspects without taking an explicitly optimal approach. That is, while resource-rational analysis could potentially explain MGV-like mechanisms as approximations to optimal behaviour, MGV cannot explain resource-rational optima as emergent from its architecture.
Yet this limitation is a consequence of methodological choice rather than oversight. We prioritised preserving psychological detail over deriving mechanisms from first principles. The result is a framework that describes what metacognitive components might look like but cannot explain why they should take those forms rather than others. Whether dual-counter FOK approximates optimal stopping behaviour, whether satisficing thresholds should decay exponentially rather than linearly, whether tripartite knowledge organisation is functionally necessary: these questions cannot be answered within MGV but could potentially be addressed through resource-rational analysis. The grounding of MGV in resource-rational principles represents one path toward resolving this limitation.
4.4 Future Directions
Grounding MGV in Resource-Rational Principles
The most productive path forward may involve selective integration of the two approaches. Some MGV constructs appear directly amenable to resource-rational analysis. The stopping dynamics formalised in Nelson and Narens’ retrieval model parallel the optimal stopping structure of memory recall in the resource-rational framework, where feeling-of-knowing serves as the monitoring signal that enables efficient termination of searches unlikely to succeed (Callaway et al., 2024). Flavell’s cognitive monitoring, which involves inferring the state of ongoing cognitive processes from imperfect signals, shares the partial observability structure of the belief-MDP formulation of cognitive control (Lieder et al., 2018). The threshold dynamics that govern when to cease retrieval or when to shift strategies may be derivable from VOC calculations rather than stipulated as free parameters.
Other constructs may serve primarily descriptive functions that complement rather than compete with normative analysis. The tripartite knowledge taxonomy (person, task, and strategy knowledge) provides a vocabulary for the features over which resource-rational agents must learn, but the taxonomy itself is not derived from optimality principles. Similarly, the phenomenological vocabulary of metacognitive experiences (feeling-of-knowing, judgment-of-learning, feeling-of-confidence) names the signals that enable metacognitive control, but resource-rational analysis explains why such signals are useful rather than what they feel like. A complete theory may require both the normative grounding that resource-rational analysis provides and the psychological granularity that MGV preserves.
Meta-Reasoning and Meta-Learning in Language Models.
Griffiths et al. (2019) identify meta-reasoning and meta-learning as two components of human intelligence that current AI systems lack: meta-reasoning concerns efficient allocation of computational resources within a task, while meta-learning concerns efficient use of data across tasks to accelerate future learning. Resource-rational analysis provides formal frameworks for both (see Appendix 7). MGV formalises both components from a psychological rather than normative starting point. For meta-reasoning, the models derived from Flavell’s cognitive monitoring and Nelson and Narens’ retrieval address how to allocate effort during task performance, specifying when the Monitor should trigger the Generator to produce new reasoning and when the Verifier should accept or reject candidate outputs. For meta-learning, the model derived from Nelson and Narens’ acquisition addresses how to distribute study across items for future retrieval, deciding which items merit additional encoding effort given current memory states.
Recent work has begun applying rational metareasoning principles to language models. De Sabbata et al. (2024) developed RaM, a VOC-inspired reward function that balances the utility of chain-of-thought reasoning against its token cost. With RaM, models learn to adapt reasoning length to task difficulty, using substantially fewer tokens on easier problems. This demonstrates that VOC-based training can induce adaptive computation, but it addresses only one dimension of meta-reasoning: how long to <think>. The resource-rational frameworks reviewed in Appendix 7 reveal additional dimensions, including what to <think> about (as in planning, where the agent selects which nodes to expand), how to <think> (as in strategy selection, where the agent chooses among qualitatively different reasoning modes), and how intensely to <think> (as in cognitive control, where the agent modulates the strength of control signals). MGV provides architectural mechanisms for these decisions. A natural direction for future work is to ground these mechanisms in resource-rational principles, deriving MGV’s control dynamics from VOC optimisation rather than stipulating them directly.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and introduction clearly state this is a theoretical contribution that formalises metacognitive theories, not an empirical study claiming performance improvements.
  5. 2.
    Limitations
  6. Question: Does the paper discuss the limitations of the work performed by the authors?
  7. Answer: [Yes]
  8. Justification: The paper explicitly states it makes no empirical claims and that translating psychological concepts to computational systems remains speculative.
  9. 3.
    Theory assumptions and proofs
  10. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  11. Answer: [N/A]
  12. Justification: This paper formalises existing psychological theories into computational frameworks rather than proving new theoretical results.
  13. 4.
    Experimental result reproducibility
  14. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results?
  15. Answer: [N/A]
  16. Justification: The paper does not include experiments; it presents theoretical formalisations of cognitive science theories.
  17. 5.
    Open access to data and code
  18. Question: Does the paper provide open access to the data and code?
  19. Answer: [N/A]
  20. Justification: The paper is purely theoretical and does not involve experiments requiring code or data.
  21. 6.
    Experimental setting/details
  22. Question: Does the paper specify all the training and test details?
  23. Answer: [N/A]
  24. Justification: The paper does not include experiments.
  25. 7.
    Experiment statistical significance
  26. Question: Does the paper report error bars suitably and correctly defined?
  27. Answer: [N/A]
  28. Justification: The paper does not include experiments.
  29. 8.
    Experiments compute resources
  30. Question: For each experiment, does the paper provide sufficient information on the computer resources?
  31. Answer: [N/A]
  32. Justification: The paper does not include experiments.
  33. 9.
    Code of ethics
  34. Question: Does the research conducted in the paper conform with the NeurIPS Code of Ethics?
  35. Answer: [Yes]
  36. Justification: The theoretical work conforms with ethical guidelines and poses no ethical concerns.
  37. 10.
    Broader impacts
  38. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts?
  39. Answer: [N/A]
  40. Justification: This is foundational theoretical work without direct societal applications or deployment considerations.
  41. 11.
    Safeguards
  42. Question: Does the paper describe safeguards for responsible release?
  43. Answer: [N/A]
  44. Justification: The paper poses no risks as it neither releases models nor datasets, only theoretical formalisations.
  45. 12.
    Licenses for existing assets
  46. Question: Are the creators or original owners of assets properly credited?
  47. Answer: [N/A]
  48. Justification: The paper does not use existing code, data, or model assets.
  49. 13.
    New assets
  50. Question: Are new assets introduced in the paper well documented?
  51. Answer: [N/A]
  52. Justification: The paper does not release new assets.
  53. 14.
    Crowdsourcing and research with human subjects
  54. Question: For crowdsourcing experiments and research with human subjects, does the paper include full details?
  55. Answer: [N/A]
  56. Justification: The paper does not involve crowdsourcing or research with human subjects.
  57. 15.
    Institutional review board (IRB) approvals
  58. Question: Does the paper describe potential risks incurred by study participants?
  59. Answer: [N/A]
  60. Justification: The paper does not involve research with human subjects.
  61. 16.
    Declaration of LLM usage
  62. Question: Does the paper describe the usage of LLMs if it is an important component of the core methods?
  63. Answer: [N/A]
  64. Justification: LLMs are not used as a component of the core theoretical formalisations presented.